OpenStudy (anonymous):

27^(2/3) explained please

OpenStudy (anonymous):

so we can do it in steps \[((27)^{\frac{1}{3}})^2\] first take cube root of 27; 3*3*3=27 so cuberoot is 3 we then square it 3^2=9

OpenStudy (anonymous):

great explanation, thank you